# DIVE: Taming DINO for Subject-Driven Video Editing

### v1

@@ Abstract
Building on the success of diffusion models in image generation and editing, video editing has recently gained substantial attention. However, maintaining temporal consistency and motion alignment still remains challenging. To address these issues, this paper proposes DINO-guided Video Editing (DIVE), a framework designed to facilitate subject-driven editing in source videos conditioned on either target text prompts or reference images with specific identities. The core of DIVE lies in leveraging the powerful semantic features extracted from a pretrained DINOv2 model as implicit correspondences to guide the editing process. Specifically, to ensure temporal motion consistency, DIVE employs DINO features to align with the motion trajectory of the source video. For precise subject editing, DIVE incorporates the DINO features of reference images into a pretrained text-to-image model to learn Low-Rank Adaptations (LoRAs), effectively registering the target subject’s identity. Extensive experiments on diverse real-world videos demonstrate that our framework can achieve high-quality editing results with robust motion consistency, highlighting the potential of DINO to contribute to video editing. Project page: https://dino-video-editing.github.io.
@@ 1 Introduction
Compared to image editing [24, 3, 48, 30], video editing [56, 70, 76, 43, 35] presents unique challenges, especially in capturing the motion from source videos and transferring it to edited results while maintaining temporal consistency. Most methods address this challenge by relying on pretrained T2I models, introducing additional temporal layers, motion modules, or unique mechanisms such as spatial-causal attention and temporal attention [65, 35, 28, 12]. These methods typically transfer the source video’s motion to the target video by injecting temporal attention maps or features. For instance, Tune-A-Video [65] performs cross frame attention between each frame and anchor frames (e.g., the first and previous frames) to preserve appearance consistency. TokenFlow [12] further explicitly propagates diffusion features based on inter-frame correspondences to improve consistency. However, the stored features in these extra modules or internal blocks may inadvertently retain some source appearance details, leading to a blend of source and target appearances and resulting in visual incoherence.
@@
In this paper, we aim to establish more robust video correspondences that not only contain minimal appearance information but also can accurately align motion trajectories without manual annotation. Motivated by recent advances in self-supervised learning, we leverage a pretrained DINOv2 model [41], a vision Transformer distilled using a large collection of natural images. DINO features have been demonstrated to have robust and discriminative capabilities in capturing fine-grained semantic information, proving valuable across visual tasks [1, 39, 50]. Additionally, Zhang et al. [73] reveal that DINO features excel at obtaining sparse and accurate matches on image semantic correspondence. However, their application in the video domain has not been fully explored. Based on these insights, we introduce DINO to the video domain and find that DINO features across video frames exhibit high semantic similarity, indicating their potential as robust video correspondences to guide the editing process.
@@
  • We demonstrate that tamed DINO features can serve as strong and robust correspondences for capturing temporal motion.
@@ 2.1 Diffusion-Based Video Generation and Editing
Diffusion models have recently emerged as a powerful approach in various vision tasks, particularly in image generation and editing [44, 74, 40, 3, 23]. Building on their success, video generation and editing with diffusion models have garnered significant research interest. Early efforts focus primarily on training video diffusion models directly in pixel or latent spaces [21, 19, 2, 61], which require extensive datasets [69, 62]. However, recent studies have shifted towards training-free models to reduce computational costs. For instance, Text2Video-Zero [26] leverages the pretrained Stable Diffusion model for video synthesis, employing the cross-attention mechanism with the first frame to maintain frame consistency.
As some focus shifts from video generation to video editing, research has increasingly explored methods that adapt pretrained T2I models, driven by their outstanding generative capabilities [65, 12, 28, 31, 60, 75, 32]. These methods often introduce additional temporal layers, motion modules, or spatial-temporal attention to achieve temporal coherence. For example, Tune-A-Video [65] incorporates a temporal self-attention layer and updates projection matrices in each attention block of a T2I model to enhance temporal modeling. FateZero [43] preserves motion and structural information by storing comprehensive attention maps at each stage of the inversion process, which are then fused during editing. Some other methods use auxiliary sequences as correspondences, such as flow [33, 6, 22], depth [9, 25, 8, 34], or edge [70, 76] maps, to model temporal motion. For instance, FLATTEN [6] integrates optical flow into diffusion models through flow-guided attention, aligning patches from different frames along consistent flow paths. However, these methods often suffer from flickering and content incoherence due to the high density of the auxiliary sequences. To address these issues, VideoSwap [15] employs semantic points as sparse video correspondences to align motion trajectories, but it requires users to manually define these points in source videos. In contrast, our approach leverages the self-supervised DINOv2 features as implicit video correspondences, enabling accurate motion alignment without the need for manual annotation.
@@ 4.1 Experimental Setup
We evaluate our method on a dataset of 30 videos curated from DAVIS [42] and the Internet, covering both objects and animals. Each video sample consists of 16 frames with a temporal stride of 4 to match the temporal window of the motion layer in AnimateDiff. We crop and reisze these videos to two alternate resolutions for editing ($H\times W$): $512\times 512$ and $512\times 896$. For videos at $512\times 512$, we further resize them to $448\times 448$ for DINO feature extraction. This preprocessing allows us to match feature scales accurately. For instance, the DINO features have a spatial dimension of $16\times 16$, while the encoder features $\bm{F}^{t}=\{\bm{F}_{1}^{t},\bm{F}_{2}^{t},\bm{F}_{3}^{t},\bm{F}_{4}^{t}\}$ of the Stable Diffusion U-Net are at scales of $\{64\times 64,32\times 32,16\times 16,8\times 8\}$. Similarly, for $512\times 896$ videos, we resize them to a $448\times 768$ for DINO feature extraction correspondingly. We further collect 8 target identity categories (4 for animals and 4 for objects), with each category containing 3 to 5 reference images curated from DreamBench [45] and the Internet. The reference images undergo the same preprocessing steps as the videos. Beyond image guidance, we also perform text-guided subject editing on the source videos for a comprehensive comparison with other methods.
In terms of implementation settings, in stage 1, we use the Adam optimizer with a learning rate of $5\times 10^{-4}$, optimizing for 50 to 100 iterations. In stage 2, we use the Adam optimizer with a learning rate of $1\times 10^{-4}$ and optimize for 800 to 1000 iterations. In stage 3, we employ 50 steps of DDIM inversion and 50 steps of denoising, with $\lambda$ in Eq. 3 set to 1.0. All experiments are conducted on a single NVIDIA RTX 3090 GPU.
@@ 4.2 Qualitative Comparison
We compare DIVE against 4 baseline methods: TokenFlow [12], AnyV2V [28], FLATTEN [6], and RAVE [25], where the former two are attention-based and the latter two are correspondence-based. For all baseline methods, we use the default settings provided in their official GitHub repositories. The comparison results on various videos with different reference images are shown in Figure 4. As observed, DIVE can accurately replace the source subject’s identity with that of the target based on the reference images, while aligning with the source motion trajectory and preserving the background context. In contrast, TokenFlow and AnyV2V struggle to maintain motion alignment and often introduce appearance details from the source subject, as simple attention transfer techniques are limited in their ability to disentangle motion and appearance effectively. Although RAVE and FLATTEN can model the source motion by relying on correspondences like optical flows and depth map sequences, they tend to produce blurred and inconsistent results due to the high density of these correspondences. Additionally, to demonstrate DIVE’s generalization to text-guided video editing, we omit the identity registration step (stage 2 in Figure 3) and directly replace the source subject with the target one in the text prompt during inference in stage 3. Visual comparisons with other methods are presented in Figure 5, where DIVE shows superior temporal consistency by aligning the source motion across all parts of the subject and achieving coherent content across frames. More results are provided in the supplementary materials.
@@ 4.3 Quantitative Comparison
For quantitative evaluation, we use the metrics proposed in [29, 66], assessing DIVE and the baseline methods across three key dimensions: (1) Text Alignment, by computing the text-image similarity in the CLIP feature space for each frame-prompt pair; (2) Image Alignment, by measuring the visual similarity between each edited frame and the reference images using CLIP image features; and (3) Temporal Consistency, by calculating the similarity of CLIP features across consecutive frames. Additionally, we conduct a user study with 50 participants, asking them to select their preferred results among the different methods. The results, summarized in Table 1, demonstrate that DIVE consistently outperforms other methods across all quantitative metrics and achieves high preference rates in the user study. This superior performance highlights DIVE’s ability to produce high-quality, temporally consistent subject edits that align well with the given text prompts or reference images.
Methods Text Alignment ()($\uparrow$)( ↑ ) Image Alignment ()($\uparrow$)( ↑ ) Temporal Consistency ()($\uparrow$)( ↑ ) User Study ()($\uparrow$)( ↑ )
@@ 4.4 Ablation Study
Motion Guidance. To validate the effectiveness of DINO features as temporal motion guidance, we remove the DINO feature extraction branch and the learnable MLPs $\bm{\psi}$ in Eq. 3, but fine-tune the motion module from AnimateDiff [16]. The editing result, shown in the second row of Figure 6, indicates that although this setup can capture the source motion, it struggles to disentangle the motion from the appearance of the source subject. In contrast, our method can fully edit the subject without introducing elements of the source appearance, as DINO features exhibit high sparsity and semantic consistency across frames (see the last row). Additionally, to further evaluate the effectiveness of DINO features in capturing source motion trajectories, we vary the guidance injection weighting parameter $\lambda$ in Eq. 3 while keeping the motion module frozen. The comparison in the last three rows of Figure 6 shows that without adequate DINO feature guidance, the motion cannot be consistently maintained, demonstrating the robustness of DINO features as video correspondences for reliable video editing.
Figure 6: Ablation studies of motion guidance in stage 1.
@@
Identity Guidance. To evaluate the impact of DINO features on identity registration, we perform an ablation study with and without DINO guidance. Without DINO guidance (second row in Figure 7), the edited corgi incorrectly has a tail, even though the reference images show a tail-less corgi. This error occurs because the model, without DINO’s semantic guidance, relies on default biases from the pretrained diffusion model. In contrast, with DINO guidance (bottom row), the edited corgi correctly matches the reference images, without a tail. This demonstrates that DINO features provide accurate, part-level guidance, allowing the model to follow the reference images closely and achieve high semantic fidelity to the target identity.
Figure 7: Ablation studies of identity guidance in stage 2.
5 Conclusion and Limitations
@@
This paper presents DIVE, a framework for subject-driven video editing that leverages DINO features to achieve reliable motion alignment and precise identity registration. DIVE effectively replaces the subject in a source video with a specified target identity, whether described by reference images or text prompts, while preserving the original motion trajectory and background context. Extensive qualitative and quantitative comparisons demonstrate that DIVE outperforms existing methods in visual quality, motion consistency, and identity alignment, highlighting the potential of DINO to advance video editing.
Despite its strengths, DIVE has certain limitations. First, it may produce artifacts in preserving the target subject’s identity and struggle with complex scenes involving intricate interactions, occlusions, or large view changes due to the limited generative capacity of the pretrained base model. Additionally, DIVE can be more time-consuming compared to zero-shot methods, as it requires individual fine-tuning for each editing case. Future work could focus on optimizing DIVE for efficiency and improved generalization.

### v2

@@ Abstract
Building on the success of diffusion models in image generation and editing, video editing has recently gained substantial attention. However, maintaining temporal consistency and motion alignment still remains challenging. To address these issues, this paper proposes DINO-guided Video Editing (DIVE), a framework designed to facilitate subject-driven editing in source videos conditioned on either target text prompts or reference images with specific identities. The core of DIVE lies in leveraging the powerful semantic features extracted from a pretrained DINOv2 model as implicit correspondences to guide the editing process. Specifically, to ensure temporal motion consistency, DIVE employs DINO features to align with the motion trajectory of the source video. For precise subject editing, DIVE incorporates the DINO features of reference images into a pretrained text-to-image model to learn Low-Rank Adaptations (LoRAs), effectively registering the target subject’s identity. Extensive experiments on diverse real-world videos demonstrate that our framework can achieve high-quality editing results with robust motion consistency, highlighting the potential of DINO to contribute to video editing.
@@ 1 Introduction
Compared to image editing [25, 3, 51, 32], video editing [59, 73, 79, 46, 37, 44] presents unique challenges, especially in capturing the motion from source videos and transferring it to edited results while maintaining temporal consistency. Most methods address this challenge by relying on pretrained T2I models, introducing additional temporal layers, motion modules, or unique mechanisms such as spatial-causal attention and temporal attention [68, 37, 30, 13]. These methods typically transfer the source video’s motion to the target video by injecting temporal attention maps or features. For instance, Tune-A-Video [68] performs cross frame attention between each frame and anchor frames (e.g., the first and previous frames) to preserve appearance consistency. TokenFlow [13] further explicitly propagates diffusion features based on inter-frame correspondences to improve consistency. However, the stored features in these extra modules or internal blocks may inadvertently retain some source appearance details, leading to a blend of source and target appearances and resulting in visual incoherence.
@@
This paper aims to establish more robust video correspondences that not only contain minimal appearance information but also can accurately align motion trajectories without manual annotation. Motivated by recent advances in self-supervised learning, we leverage a pretrained DINOv2 model [43], a vision Transformer distilled using a large collection of natural images. DINO features have been validated to be robust and discriminative in capturing fine-grained semantic information, proving valuable across visual tasks [1, 41, 53]. Additionally, Zhang et al. [76] reveal that DINO features excel at obtaining sparse and accurate matches on image semantic correspondence. However, their application in the video domain has not been fully explored. Based on these insights, we introduce DINO to the video domain and find that DINO features across video frames exhibit high semantic similarity, indicating their potential as robust video correspondences to guide the editing process.
@@
  • We show that tamed DINO features offer strong and robust correspondences for capturing temporal motion.
@@ 2.1 Diffusion-Based Video Generation and Editing
Diffusion models have recently emerged in various vision tasks, especially in image generation and editing [47, 77, 42, 3, 24]. Building on their success, video generation and editing with diffusion models have attracted significant research interest. Early efforts focus primarily on training video diffusion models directly in pixel or latent spaces [22, 20, 2, 64], which require large datasets [72, 65]. However, recent studies have shifted towards training-free models to reduce computational costs. For instance, Text2Video-Zero [28] leverages the pretrained Stable Diffusion model for video synthesis, employing cross-attention with the first frame to maintain frame consistency.
As focus shifts from video generation to editing, research has increasingly explored methods that adapt pretrained T2I models, driven by their strong generative capabilities [68, 13, 30, 33, 63, 78, 34, 6]. These methods often introduce extra temporal layers, motion modules, or spatial-temporal attention to ensure temporal coherence. For example, Tune-A-Video [68] adds temporal self-attention and updates projection matrices in each attention block to enhance temporal modeling. FateZero [46] stores attention maps at each inversion stage and fuses them during editing to preserve motion and structure. Others rely on auxiliary sequences, such as flow [35, 7, 23], depth [10, 27, 9, 36], or edge [73, 79] maps, to model motion. FLATTEN [7], for instance, uses flow-guided attention to align patches across frames. However, these dense signals often cause flickering and incoherence. VideoSwap [16] mitigates this with sparse semantic points as correspondences, but it requires manual point definition. In contrast, we leverage self-supervised DINOv2 features as implicit correspondences, enabling accurate motion alignment without manual annotation.
@@ 4.1 Experimental Setup
[…]
In terms of implementation settings, in stage 1, we use the Adam optimizer with a learning rate of $5\times 10^{-4}$, optimizing for 50 to 100 iterations. In stage 2, we use the Adam optimizer with a learning rate of $1\times 10^{-4}$ and optimize for 800 to 1000 iterations. In stage 3, we employ 50 steps of DDIM inversion and 50 steps of denoising, with $\lambda$ in Eq. 3 set to 1.0. All experiments are conducted on one single NVIDIA RTX 3090 GPU.
@@ 4.2 Qualitative Comparison
We compare DIVE against 4 baseline methods: Slicedit [6], AnyV2V [30], FLATTEN [7], and RAVE [27], where the former two are attention-based and the latter two are correspondence-based. The comparison results on various videos with different reference images are shown in Figure 4. As observed, DIVE can accurately replaces the source subject’s identity with that of the target based on reference images while aligning with the source motion trajectory and preserving the background. In contrast, Slicedit and AnyV2V struggle to maintain motion alignment and often introduce appearance details from the source subject, as simple attention transfer techniques are limited in disentangling motion and appearance effectively. Although RAVE and FLATTEN model the source motion using correspondences like optical flows and depth maps, they tend to produce blurred and inconsistent results due to the high density of these correspondences.
Additionally, to demonstrate DIVE’s generalization to text-guided video editing, we omit the identity registration step (stage 2 in Figure 3) and directly replace the source subject with the target one in the text prompt during inference in stage 3. Visual comparisons with other methods are presented in Figure 5, where DIVE shows superior temporal consistency by aligning the source motion across all parts of the subject and achieving coherent content across frames. More results are provided in the supplementary materials.
@@ 4.3 Quantitative Comparison
For quantitative evaluation, we use several metrics to assess DIVE and the baseline methods, which are (1) Text Alignment [69], by computing the text-image similarity in the CLIP feature space for each frame-prompt pair; (2) Image Alignment [31], by measuring the visual similarity between each edited frame and the reference images using CLIP image features; (3) Temporal Consistency [69], by calculating the similarity of CLIP features across consecutive frames; and (4) Overall Video Quality [26], by calculating scores from multiple dimensions and averaging them. Additionally, we conduct a user study with 50 participants, asking them to select their preferred results among these methods. The results listed in Table 1, demonstrate that DIVE consistently outperforms other methods across all quantitative metrics and achieves high preference rates in the user study. This superior performance highlights DIVE’s ability to produce high-quality, temporally consistent subject edits that align well with the given text prompts or reference images.
Method
Text
Alignment [69] $\uparrow$
Image
Alignment [31] $\uparrow$
Temporal
Consistency [69] $\uparrow$
Overall
Video Quality [26] $\uparrow$
User
Study $\uparrow$
@@ 4.4 Ablation Study
Motion Guidance. To validate the effectiveness of DINO features as temporal motion guidance in stage 1, we remove the DINO feature extraction branch and the learnable MLPs $\bm{\psi}$ in Eq. 3, but fine-tune the motion module from AnimateDiff [17]. The editing result shown in the second row of Figure 6 and Table 2 indicates that although this setup can capture the source motion, it struggles to disentangle the motion from the appearance of the source subject, leading to image misalignment. In contrast, our method can fully edit the subject without introducing elements of the source appearance, as DINO features exhibit high sparsity and semantic consistency across frames (see the last row). Additionally, to further evaluate the effectiveness of DINO features in capturing source motion trajectories, we vary the guidance injection weighting parameter $\lambda$ in Eq. 3 while keeping the motion module frozen. The comparison in Table 2 and the last three rows of Figure 6 show that without adequate DINO feature guidance, the motion cannot be consistently maintained, demonstrating the robustness of DINO features as video correspondences for reliable video editing.
@@
Identity Guidance. To evaluate the impact of DINO features on identity registration in stage 2, we perform an ablation study with and without DINO guidance. Results are shown in Figure 7 and Table 2. Without DINO guidance, subjects in edited videos tend to be less aligned with reference images. For example, the edited corgi incorrectly has a tail, though the reference images show a tail-less corgi. This error occurs because, without DINO’s semantic guidance, the model relies on biases from the pretrained diffusion model. In contrast, with DINO guidance, the edited corgi correctly matches the reference images. This demonstrates that DINO features provide accurate, part-level guidance, allowing the model to follow reference images closely and achieve high semantic fidelity to the target identity.
Figure 6: Visual ablations of motion guidance in stage 1.
Figure 7: Visual ablations of identity guidance in stage 2.
Implementation
Text
Alignment [69] $\uparrow$
Image
Alignment [31] $\uparrow$
Temporal
Consistency [69] $\uparrow$
Overall
Video Quality [26] $\uparrow$
Motion Guidance
w/o DINO, 29.91 67.49 92.18 0.737
w/ learnable motion module
$\lambda=0$, 28.78 79.50 87.24 0.581
w/ frozen motion module
$\lambda=0.3$, 29.05 80.28 89.53 0.628
w/ frozen motion module
$\lambda=1.0$, 29.43 84.27 92.33 0.775
w/ frozen motion module
Identity Guidance
w/o DINO 29.25 77.36 92.12 0.722
w/ DINO 29.43 84.27 92.33 0.775
Table 2: Quantitative ablations of motion and identity guidance.
5 Conclusion
@@
This paper presents DIVE, a framework for subject-driven video editing that leverages DINO features as robust video correspondences to achieve reliable motion alignment and precise identity registration. DIVE effectively replaces the subject in a source video with a specified target identity, whether described by reference images or text prompts, while preserving the original motion trajectory and background context. Extensive qualitative and quantitative comparisons demonstrate that DIVE outperforms state-of-the-art methods in visual quality, motion consistency, and identity alignment, highlighting the potential of DINO to contribute to video editing.
Acknowledgement
This work was supported by the Shenzhen Science and Technology Program (JSGG20220831105002004) and Shenzhen Key Laboratory of Computer Vision and Pattern Recognition.